# appAppendix References

### v5

@@ Abstract
As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization has attracted growing attention. Existing works in this field aim to establish an understanding of whether or to what extent DPMs learn via memorization. Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI. Existing works revealed that conditional DPMs are more prone to memorize training data than unconditional DPMs. And most data extraction methods developed so far target conditional DPMs. Although unconditional DPMs are less prone to data extraction, further investigation into these attacks remains essential since they serve as the foundation for conditional models like Stable Diffusion, and exploring these attacks will enhance our understanding of memorization in DPMs. In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on generated data as surrogate conditions to extract training data from unconditional DPMs. Empirical results demonstrate that it can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset. Furthermore, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs and why SIDE is effective.
@@ 1 Introduction
The diffusion probabilistic models (DPMs) [16, 30, 34] is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion [26], DALL-E 3 [2], and Sora [3]. Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches or copyright infringement, as highlighted in the literature [31, 32, 1, 11]. A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training data[4]. Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by [5] demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion [26], revealing the potential privacy and copyright dangers associated with them.
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., conditional-trained DPMs are more prone to memorization [32]. Gu et al. [13] investigates influential factors on memorization behaviors via comprehensive experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though these works indicate that unconditional diffusion models are much less prone to data extraction attacks, investigating such attacks for unconditional diffusion models is still necessary. This is because 1) unconditional DPMs are the foundations of conditional generative models like Stable Diffusion, which adapts pre-trained unconditional DPMs as the essential part for conditional generation or combines them with conditional training approaches [26, 39]; 2) exploring data extraction attacks on unconditional DPMs enhances understanding of memorization mechanisms in both conditional and unconditional DPMs.
@@
In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs. This attack method uses a classifier trained on the same or similar training data as a surrogate condition for unconditional DPMs. SIDE can extract training data effectively by making the classifier time-dependent (to the diffusion sampling process). We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1). Additionally, to offer a theoretical analysis, we introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner. Based on this metric, we explain why conditional DPMs are more likely to memorize training data, why random labelling can increase this memorization and why the SIDE method is effective for extracting training data in unconditional DPMs.
@@
  • We propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical analysis that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) why the SIDE method is effective for extracting training data in unconditional DPMs.
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets and show that, on average, it can outperform the baseline method by more than 50%.
Figure 1: A few examples of the extracted images from a DDPM trained on a subset of the CelebA dataset using our SIDE method. Top: training images; bottom: extracted images.
@@ 2 Related Work
Diffusion Probabilistic Models  DPMs [30] such as Stable Diffusion [26], DALL-E 3 [2], Sora [3], Runway [26], and Imagen [28] have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks [12]. These models can be viewed from two perspectives. The first is score matching [34], where diffusion models learn the gradient of the image distribution [35]. The second perspective involves denoising DPMs [16], which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, [12] utilizes a classifier to guide the denoising process at each sampling step. Additionally, [17] introduces classifier-free guidance for conditional data sampling using DPMs.
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models [6, 19], which has motivated more in-depth research on DPMs [32, 13, 24]. Notably, a recent research on image-generation DPMs: [31] found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by [5] in broader experiments for both conditional and unconditional diffusion models. Further studies [32] and [13] linked memorization to model conditioning, showing that conditional models are more prone to memorization. Several works have proposed methods to detect and mitigate memorization to address this issue. Wen et al. [38] proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions. It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization. Ren et al. [25] proposed two metrics based on findings on cross-attention in DPMs to detect memorization. To mitigate memorization, Chen et al.[7] proposed anti-memorization guidance to mitigate memorization during the sampling process. Instead of using guidance, Ren et al. [25] mitigated memorization by changing the attention score or masking out summary tokens in the cross-attention layer. Wen et al.[38] achieved this by minimizing the magnitude during inference. So far, the motivated data extraction attack or defend methods are primarily for conditional DPMs [5, 37], and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs [13]. Though challenging, investigating data extraction methods in unconditional DPMs is still necessary. This is because 1) Unconditional DPMs are key parts in conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) Exploring this approach enhances our understanding of memorization in DPMs. In response to these challenges, we propose a novel data extraction method called SIDE for unconditional DPMs and providing theoretical understanding of SIDE.
3 Proposed SIDE Method
3.1 Intuition
@@
Previous studies [13, 32] have shown that conditional DPMs are more prone to memorization due to their reliance on explicit conditioning labels, such as class labels or text conditions, which we refer to as explicit informative labels or explicit labels. In contrast, unconditional DPMs do not have access to these explicit labels. Nevertheless, they still learn to form representation clusters from the training data, even though these clusters are not directly specified in the data itself [8]. We name these representation clusters as implicit informative labels or implicit labels. By formalizing the clustering patterns within the training data, it may be possible to construct implicit labels and make the unconditional DPMs conditional on these implicit labels to extract training data from unconditional DPMs.
In the following section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
3.2 Constructing Implicit Informative Labels
@@
Intuitively, one could use a classifier to generate (predict) implicit labels $y_{I}$ during the sampling process of the diffusion model. The classifier can be a normal classifier trained on the same dataset as the diffusion model. When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels. We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process:
@@
$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (1)
where $x$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
We can use the classifier that generates the implicit labels to approximate the gradient in Eq. (3.2). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration [14]. Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right)$ (2)
Then, we have:
@@
$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (3)
[…]
Figure 2: An illustration of our proposed Time-Dependent Knowledge Distillation (TDKD) method that trains a time-dependent classifier on a pseudo-labeled synthetic dataset.
3.3 Time-dependent Classifier
In Eq. (3.2), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to distil knowledge from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques [10, 9]. We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as follows:
@@ 3.3 Time-dependent Classifier
Overall Procedure of SIDE

### v6

@@ Abstract
As diffusion probabilistic models (DPMs) become the mainstream framework for Generative AI (GenAI), understanding their memorization behavior is critical for assessing risks of data leakage, copyright infringement, and ensuring trustworthy deployment. Previous studies indicate that conditional DPMs are more prone to memorization compared to unconditional DPMs, with the majority of extraction techniques focusing on conditional models. However, studying unconditional DPMs remains crucial, as they serve as the foundation for advanced models such as Stable Diffusion. Understanding their memorization behavior can provide valuable insights into the broader characteristics of DPMs. In this work, we propose Surrogate condItional Data Extraction (SIDE), a novel data extraction method for unconditional DPMs. SIDE employs a time-dependent classifier trained on generated data as surrogate conditions to extract training samples. Empirical results show that SIDE extracts training data effectively in challenging scenarios where existing methods fail, with an average performance gain of over 50% on the CelebA dataset across different scales. We also provide a unified theoretical framework explaining memorization in both conditional and unconditional DPMs and why SIDE works. Our findings contribute to a deeper understanding of memorization in GenAI and provide practical guidance for mitigating associated risks.
@@ 1 Introduction
Diffusion probabilistic models (DPMs) [21, 38, 42] are a family of powerful generative models that learn data distributions by progressively corrupting data through a forward diffusion process and then reconstructing it via a reverse diffusion process. Due to their exceptional capability to capture complex data distributions, DPMs have become foundation models for many pioneering Generative Artificial Intelligence (GenAI) products, including Stable Diffusion [33], DALL-E 3 [4], and Sora [5].
However, the growing adoption of DPMs raises concerns about data memorization — the risk of memorizing raw training samples. This may result in the generation of duplicated rather than novel content, increasing the risk of data leakage, privacy violations, and copyright infringement. [39, 40, 2, 13]. For instance, Stable Diffusion has been criticized as a “21st-century collage tool” that remixes copyrighted works from artists whose data was used for training [6]. Moreover, data memorization enables data extraction attacks, a type of privacy attack aimed at recovering training data from trained models. Recent work by [7] demonstrated the feasibility of extracting training data from DPMs like Stable Diffusion [33], highlighting significant privacy and copyright risks.
Existing studies have explored data memorization in DPMs, revealing that conditional DPMs exhibit a higher tendency to memorize training data than unconditional DPMs [40]. For example, [18] found that conditioning on random-labeled data significantly increases memorization, while unconditional models exhibit much lower memorization rates. Despite these findings, investigating data extraction attacks on unconditional DPMs remains crucial for two reasons: (1) unconditional DPMs form the foundation of conditional models like Stable Diffusion, which adapt pre-trained unconditional DPMs for conditional generation [33, 48]; and (2) understanding memorization in unconditional DPMs provides insights into the broader memorization mechanisms of DPMs.
@@
In this work, we propose Surrogate condItional Data Extraction (SIDE), a novel method for extracting training data from unconditional DPMs. SIDE employs a time-dependent classifier trained on generated data as surrogate conditions to effectively extract training samples. We empirically validate SIDE on CIFAR-10, multiple scales of the CelebA and ImageNet, demonstrating its superiority over baseline methods by achieving over 50% higher effectiveness on average. Examples of extracted images are shown in Figure 1. Additionally, we introduce a divergence to measure memorization in DPMs and present a theoretical analysis explaining: (1) why conditional DPMs memorize more data (even labelled with random labels), (2) why SIDE is effective for extracting training data.
@@
  • We propose SIDE, a novel method that leverages implicit labels provided by a time-dependent classifier to extract training data from unconditional DPMs.
  • We introduce a memorization divergence measure and provide a theoretical analysis that explains the impact of conditioning in DPMs and the effectiveness of SIDE.
  • We empirically validate the effectiveness of SIDE on CIFAR-10, CelebA and ImageNet, demonstrating it outperforms baseline methods by over 50% on average.
Figure 1: Examples of training images (top) and corresponding extracted images (bottom) from a DDPM trained on a subset of CelebA using our SIDE method.
@@ 2 Related Work
DPMs [38] have achieved state-of-the-art performance in image and video generation, as exemplified by models such as Stable Diffusion [33], DALL-E 3 [4], Sora [5], Runway [33] and Imagen [35]. These models excel on various benchmarks [15]. DPMs can be interpreted from two perspectives: 1) score matching [42], where model learns the gradient of data distribution [43], and (2) denoising diffusion [21], where Gaussian noise is added to clean images over multiple time steps, and the model is trained to reverse this process. For conditional sampling, [15] introduced classifier guidance to steer the denoising process, while [22] proposed classifier-free guidance, enabling conditional generation without explicit classifiers.
Memorization in Diffusion Models
Early research on memorization primarily focused on language models [8, 25], which later inspired subsequent studies on DPMs [40, 18, 31, 36, 16, 1, 16, 27, 47, 3, 17]. Notably, [39] found that 0.5-2% of generated images duplicate training samples, a result corroborated by [7] through more extensive experiments on both conditional and unconditional DPMs. Further studies [40, 18] linked memorization to model conditioning, showing that conditional DPMs are more prone to memorization. To address memorization, several methods have been proposed for detection and mitigation. For example, [46] introduced a method to detect memorization-triggering prompts by analyzing the magnitude of text-conditional predictions, achieving high accuracy with minimal computational overhead. [32] proposed metrics based on cross-attention patterns in DPMs to identify memorization. On the mitigation side, [9] developed anti-memorization guidance to reduce memorization during sampling, while [32] modified attention scores or masked summary tokens in the cross-attention layer. [46] minimized memorization by controlling prediction magnitudes during inference.
Despite these advances, the current research mostly focuses on conditional DPMs [7, 45], as extracting data from unconditional DPMs has been found to be extremely more challenging [18]. However, investigating unconditional DPMs remains critical because 1) they form the foundation of conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) understanding memorization in unconditional DPMs provides insights into the broader behavior of DPMs. In response to these challenges, we propose SIDE, a novel data extraction method for unconditional DPMs, and provide a unified theoretical framework to explain memorization in both conditional and unconditional DPMs.
3 Surrogate Conditional Data Extraction
Threat Model
In this work, we focus on unconditional DPMs and assume that the attacker has access to the target DPM as well as a classifier trained on the same or similar dataset. The attacker’s objective is to extract the original training samples from the target DPM. In cases where the classifier is unavailable, the attacker can leverage generated data to train one, following a standard training procedure, potentially with certain necessary annotations. It is important to note that the classifier does not leak training data, as we will empirically demonstrate in Sec. 5.3. Its sole purpose, in this context, is to generate implicit labels, which can alternatively be achieved through clustering techniques.
3.1 Intuition of SIDE
@@
Previous research [18, 40] have shown that conditional DPMs are more susceptible to memorization due to their reliance on explicit labels (e.g., class labels or text conditions). In contrast, unconditional DPMs lack access to these explicit labels. However, they still learn to form representation clusters from the training data, even though these clusters are not explicitly defined [10]. We term these representation clusters as implicit labels. By formalizing the clustering patterns within the training data, it becomes possible to construct implicit labels and condition unconditional DPMs on these labels to facilitate training data extraction.
In the following, we detail the construction of implicit labels for unconditional DPMs, their conversion into explicit labels, and the subsequent use of these explicit labels for training data extraction.
3.2 Constructing Implicit Labels
@@
Intuitively, implicit labels can be generated during the sampling process of the diffusion model using a classifier. This classifier can be trained on the same/similar dataset as the diffusion model. When such a classifier is unavailable, representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can serve as implicit labels. We assume that an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process described as:
@@
$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w,$ (1)
where ${\bm{x}}$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
To approximate the gradient in the sampling process, we use the classifier that generates the implicit labels. However, neural network classifiers are prone to miscalibration [19], which can affect the sampling process. To address this, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output using a power prior:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right).$ (2)
This leads to a classifier-conditioned sampling process:
@@
$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w.$ (3)
This approach differs from [15], which introduced classifier guidance for improving sample quality under the assumption that with $Z$ being a constant. However, this assumption only holds when $\lambda=1$, as $Z$ is explicitly dependent on the ${\bm{x}}_{t}$ (the $t$-th step of sampling image ${\bm{x}}$ ) when $\lambda\neq 1$. Our derivation solves this issue by redefining $p^{t}_{\theta}\left(y|{\bm{x}}\right)$ using the power prior.
Figure 2: An illustration of our proposed time-dependent classifier training method on a pseudo-labeled synthetic dataset.
3.3 Training Time-dependent Classifier
In Eq. (3.2) above, the classifier is represented as $p_{\theta}^{t}(y|{\bm{x}})$, where $t$ denotes the timestep, indicating its time-dependent nature. However, we only have access to a time-dependent classifier. To address this limitation, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier based on synthetic data generated at different sampling steps. The distillation process is illustrated in Figure 2. TDKD enables the classifier to provide time-dependent guidance during sampling through a two-step process:
Step 1: The network architecture is modified to incorporate time-dependent inputs. The structure of the time-dependent module and the specific modifications are detailed in Appendix B.
Step 2: A synthetic dataset with pseudo labels is generated to distill knowledge from the time-independent classifier to its time-dependent counterpart.
At Step 2, we can leverage the target DPM to generate a synthetic dataset, following established generative data augmentation techniques [12, 11]. Pseudo labels for the synthetic dataset are then generated using the time-independent classifier trained on the original dataset. Finally, the time-dependent classifier is trained on the labeled synthetic dataset. The training objective is defined as:
@@ 3.3 Training Time-dependent Classifier
3.4 Overall Procedure of SIDE